### v1

@@ A.1 Fully convolutional network specificities
In Sects. 2.1 and 3.6, we stated that our network architecture creates a mapping from a 2D input image to a regular output grid corresponding to independent regions of the input image. We also exposed that a succession of convolution operations is appropriate for this mapping as they apply an identical filter to various regions of an image. A convolution is translational equivariant, meaning that if the input content is shifted, the convolution response is equally shifted in the same direction. Interestingly, in the traditional YOLO-V1 implementation, the last few layers were not convolutional. It means that the detection units from each grid cell had their own set of weights and that their prediction was position-dependent in the image. In this version, the box position was not structurally tied to the input area corresponding to the grid cell, and the receptive field of the network was always equal to the full image. However, this approach does not fully exploit the translational equivariance of the convolution operation to construct a real position-agnostic object detector.
@@
Starting with YOLO-V2, they switched to fully convolutional formalism, which we also adopted for our method. This type of architecture naturally reduces the spatial dimension up to the desired output grid resolution. This presents two main advantages. First, each grid cell is the direct spatial reduction of an input region of the size of the network reduction factor. For example, a network that contains four $2{\times}2$ pooling layers has a reduction factor of $162^{4}=$ 16 in both dimensions, which would map an input image of size $128{\times}128$ to an $8{\times}8$ output grid, with each element directly mapping a $16{\times}16$ region in the input image. Adding convolution layers with overlapping filters between the spatial reductions increases the input receptive field. This way, each output grid cell can search objects in a much larger area but remains ”structurally” centered on the appropriate input region. Second, a fully convolutional architecture preserves translational equivariance for shift values that are multiples of the network reduction factor. Using a convolutional layer as output, the grid cells share the same weight vector. It means that the output layer defines a single ”detector” that is applied independently to sub-parts of the image to construct the grided output. This property ensures that the objectness score has the same meaning for a given detection unit over all the grid cells for post-prediction filtering (Sect. 2.8). It also strongly helps the training process since objects centered on different grid cells contribute to training the same position-agnostic set of weights.
@@
An interesting side effect of having a fully convolutional structure is that we can use any input size that is a multiple of the reduction factor. Doing so will only change the size of the output grid but not the detector’s behavior at the level of a detection unit. In classical computer vision applications, the detections must be invariant to the apparent scale of an object and its resolutions. To do so, the network must be trained using various resolutions and encode the bounding boxes in a resolution invariant format, for example, predicting the box sizes as fractions of the image size. In the classical YOLO implementation (starting with V2), this principle is mostly used to adjust the computational cost and the detection accuracy of the same model by adjusting the input resolution used at prediction time. While this approach helps preserve resolution invariance, it is not quite suited to our astronomical applications. In our datasets, the pixel size is valuable information regarding the instrument limitations. In particular, due to the synthesized beam effect acting as a resolution limit, downscaling an image by a factor of two will not result in sources that look similar to sources of half the size. It goes the same way for the flux value that can not be trivially preserved when changing the scaling of an astronomical image. For these reasons, we decided to drop the detector invariance to the apparent scale. Instead, we consider all the predicted box quantities directly related to a fixed pixel scale constrained during training. By doing this, our detector will predict the same detection box for a specific source regardless of the input size as long as the necessary information is contained in the input field. With this, we could technically apply our detector to the full $32768{\times}32768$ input image of the SDC1 in a single prediction path. This should produce the same results as with our overlapping patches split, or even better results as it would eliminate edge effects created by the patch decomposition. We use this principle in Sects 3.8 and 3.9, independently optimizing the training and prediction input sizes regarding their respective constraints.
@@ A.2 Grid resolution and detection unit density
With a fully convolutional architecture, the output grid resolution depends on the input size and the chosen network backbone reduction factor. Because the number of detection units per grid cell is fixed before training (Sect. 2.5), the total number of detectable objects in an image is defined by the output grid resolution. This can be seen as a limit in the classical YOLO implementation since the detector is expected to be scale and resolution-invariant, which cannot be the case as it changes the number of available detection units. With our non-scale-invariant approach, the only quantity that matters is the density of objects at a given resolution. We consider our pixel size representing a fixed angular scale defined by the instrument resolution limit. From there, increasing the input size is simply equivalent to having a larger field of view in the sky, and the pixel size of the objects does not change. From this, we need to define our reduction factor and the number of detection units per grid cell to be representative of the source density in the field.
@@
In the classical YOLO method, the multiple detection units per grid cell were not designed to handle simultaneous multiple detections. In everyday-life images, objects are considered to obstruct each other. In most cases, only a single object of a given scale can be detected by a grid cell simultaneously. The multiple detection units per grid cell are just a way to distribute the objects by apparent size or aspect ratio. Due to this construction, the default YOLO association process during the training phase was not designed to detect multiple objects of similar size and centered in the same grid cell, as we discuss in Sect. A.3.
@@
From this, the classical answer to a high object density in the images is to reduce the input size represented by each grid cell so they detect mostly one object each. The issue with this approach is that it forces the network to have a higher resolution while still reaching a given expressivity, having a large enough receptive field, and being capable of constructing complex patterns. In practice, typical 2D astronomical images with high sensitivity are so crowded that reaching the appropriate reduction factor is unrealistic. This would result in a network that is computationally costly and challenging to train. While using a finner output grid resolution is helpful up to some point, we need to reach the target object density by being capable of detecting multiple similar objects in each grid cell. This can be done by switching to a prediction-aware association function as exposed in Sects. 2.6 and A.3. For our reference network backbone, we chose a reduction factor of 16 (Sect. 3.6), which is lower than the typical reduction factor of 32 from the YOLO-V2 architecture (darknet-19) but still high enough so our network remains trainable and preserves good computing performances.
@@ A.3 Prediction-aware association
To train our detector, we must be capable of associating a target box with the detection unit that will be tasked to represent it (Sect. 2.6). Due to the output grid formalism, a target box is automatically associated with the grid cell that contains its center. The role of the association function is then to associate the target box with one of the available detection units in that grid cell. In the classical YOLO implementation, the association process is designed to base its decision only on the size priors of each detection unit. For this, it will consider that all target and predicted boxes are centered on the same coordinates and only compare their respective IoU, as illustrated in Fig. 23. This association function searches for the size-prior that should theoretically be the best suited if the detection was perfectly centered. As stated in Sect. A.2, this comes from the idea that all detection units are supposed to represent mutually exclusive size regimes that are unlikely to be required simultaneously.
Figure 23: Illustration of a zero-centered comparison between a target in dashed green and different theoretical detection units in shades of red. This represents the classical YOLO association process in a grid cell.
To detect multiple small objects in a single grid cell, we need to have multiple detection units with comparable size priors. For this, we base our association function on the current position and size of the predicted box instead of their size priors. In such a case, the best theoretical associate would be counter-productive and force the network to make irrelevant associations. To illustrate this effect, we present in Fig. 24 a case with three target boxes and four predictions from independent detection units, all in the same grid cell, and the resulting association based on both the theoretical-prior association and the best current prediction association. We observe that the first one considers that no prediction is properly placed but that their size is relatively good, while the second one considers that the network is already relatively good at detecting the objects and that it is mostly the predicted size that needs to be adjusted.
Figure 24: Comparison of the target-prediction association in a grid cell using two different association functions. The color indicates which prediction was associated with each target. The association results are illustrated for the default YOLO association function that is based on the best theoretical priors (left) and for an association function that uses the best current network prediction (right).
@@ A.4 Box matching metric
Figure 25: Comparison of three IoU-based metrics in different contexts. For all cases, the reference box is the dashed green box, and two test boxes in blue and orange are compared. The IoU-based metrics are always obtained by computing between a given case and the reference.
A prediction-aware association function has to be capable of selecting which prediction is the best current representation of the target. It means that we need a box-matching metric that is capable of discriminating between all the possible predictions in our specific context. The classical IoU matching criteria defined in Sect. 2.1 with Eq. 9 present strong limitations. To illustrate them, we compare IoU computations between a reference box and two possible test boxes in different contexts in Fig. 25. The main problem is that the IoU is equal to zero for all couple of boxes that have no intersection regardless of their distance (Fig. 25, frame A). For our specific astronomical applications where objects can be much smaller than the input size mapped by a grid cell, we need to be capable of attributing a different match score to non-intersecting boxes based on their distance to the reference. For this, we can use the Generalized IoU (GIoU, Rezatofighi et al. 2019) that makes use of the smallest box C that encloses the two boxes A and B to be compared by refining the default IoU following
@@
The GIoU can take values from -1 to 1 while still being equal to the IoU in cases where C is equal to one of the two boxes. While leveraging the first limitation, the GIoU remains insufficient to distinguish between test boxes fully contained inside the reference box. In such a context, we would like to attribute a different match score based on the distance between the box centers (Fig. 25, frame B). For this, we can use the Distance IoU (DIoU, Zheng et al. 2019). This matching metric combines the classical IoU with the ratio between the distance of the box centers $A_{c}$ and $B_{c}$ and the smallest enclosing box diagonal-length $d$ in the form
@@
where $\rho$ represents the Euclidean distance between the two center coordinates. The DIoU can take values from -1 to 1 while being equal to the IoU and GIoU in cases where the two boxes have the same center, corresponding to $\rho^{2}(A_{c},B_{c})=0$ (Fig. 25, frame C and D). This last metric is well suited for astronomical applications, but it remains limited in cases where the aspect ratio of the boxes matters. For example, all the metrics we described would give the same results for two test boxes with the same center and identical surface but different aspect ratios (Fig. 25, frame D). This could be included in a more refined matching metric like the Complete IoU (CIoU, Zheng et al. 2019).
We stress that all these IoU variants are only used as match criteria. The bounding-box-related loss remains a simple sum-of-square error on the box position and size following what was presented in Sect. 2.1 and Eqs. 6 and 7. This contrasts with what is described in the papers describing these metrics, where they are presented as alternative loss functions for object detection methods. In practice, all these matching metrics are implemented within YOLO-CIANNA and can be selected regarding the application. The predicted objectness uses the corresponding metric (re-scaled between 0 and 1 if necessary) to follow Eqs. 10 and 11, so the predicted score reflects all the related subtleties.
@@ A.5 Association ordering effect
With multiple target objects and multiple predictions in the same grid cell, the association ordering starts to matter. In the original YOLO, it is done by looping over the target list for each grid cell. Having multiple targets on the same grid element would mean that they are very different in size, so they are considered unlikely to require the same detection unit. With an association process based on the current network prediction, any detection unit can be the best representation for any target regardless of its size prior. We defined our YOLO-CIANNA association function to be a search for the highest value in a matching score matrix that contains the ${\rm fIoU}$ for all the target-prediction pairs as described in Sect. 2.6. We illustrate the effect of such an association function compared to a loop over the targets and a loop over the predictions in Fig. 26. With our approach, if there are not enough detection units to accommodate all the targets, the well-detected objects will get associated first, which is not the case with the two approaches based on loops.
Interestingly, in the YOLO-V1 formalism, the output can be seen as a single grid cell of the size of the input image. This is due to the dense output layer and to the encoding of the predicted box center coordinates in a way that allows them to be positioned anywhere in the image regardless of the grid cell that predicts them. Therefore, all the association problems we identified when handling multiple objects per grid cell were also present in this V1 implementation but at the scale of the full image. We observe that several tricks similar to ours were used to mitigate the identified association issues. In particular, the association process at that time was based on the current predictions of the network instead of the best theoretical association. In a way, adopting a positioning system that forces each box to belong to a small grid element can be seen as a structural solution to the association issues, allowing the classical YOLO to switch to a non-prediction-aware association. Still, this solution alone is not viable for very small objects, as discussed in Sect. A.2.
Figure 26: Comparison of the target-prediction association in a grid cell regarding the search order. The color indicates which prediction was associated with each target. All predictions and targets have the same size and shape, respectively. The prediction-target pairs are compared using a scaled $({\rm DIoU}+1){\times}0.5{\times}100$ and the results for all pairs are displayed in a matching matrix. For each step of the association process, the rows and columns corresponding to the targets and predictions are highlighted in light red, with the intersection being the match score for this pair. For all subsequent steps, the now unavailable elements in the matrix are shown in gray. The three frames compare different association results, along with the corresponding association steps highlighted in the match score matrix, when looping over the targets (left), looping over the predictions (middle), or searching for the maximum value in the score matrix (right).
@@ A.6 Association refinements
Using a prediction-aware association implies that the network is free to try to distribute the diversity of objects over the detection units independently of their size prior. This added degree of freedom increases the training difficulty. Therefore, completely removing the theoretical box-prior association process is usually inefficient. This section describes several refinements we added to our default prediction-aware association process described in Sect. 2.6 to constrain the accessible solutions and guide the training process. Regarding our complete loss function (Eq. 14), these refinements will only affect the content of the $\mathds{1}^{match}$ and $\mathds{1}^{noobj}$ masks. Most of them occur after a target-prediction couple has been selected in the association search and are applied sequentially in a mutually exclusive way, following the blue B.2 frame in Fig. 4. Consequently, the association ordering matters, and a detection unit associated with a given target is not eligible as a possible solution for all the following targets, even during the refinement conditions. It also means that the current target-prediction pair found by the default association search is preserved if no box-prior that respects the different refinement conditions is available.
We emphasize that we describe all the available refinements in our method, whether or not they have been employed in the present study. Since most of them have a regularization effect, they can be useful to stabilize the training process while lowering the best achievable score. However, listing all the accessible refinement in the same place will allow future reference, and it should also allow other research teams to explore alternative efficient configurations that we might have missed.

### v2

@@ A.1 Fully convolutional network specificities
In Sects. 2.1 and 3.6, we stated that we use a fully convolutional to create a mapping from a 2D input image to a regular output grid corresponding to independent regions of the input image. This is the choice of most classical object detection methods due to its translation equivariance property. It presents two main advantages. Firstly, each grid cell is the direct spatial reduction of an input region of the size of the network reduction factor. For example, a network that contains four $2{\times}2$ pooling layers has a reduction factor of $162^{4}=$ 16 in both dimensions. This structure would take an input image of size $128{\times}128$ to produce a $8{\times}8$ output grid, with each element directly mapping a $16{\times}16$ input region. Adding convolution layers with overlapping filters between the spatial reductions increases the input receptive field. This way, each output grid cell can search for objects in a much larger area but remains “structurally” centered on the appropriate input region. Secondly, a fully convolutional architecture preserves translational equivariance for input shift values that are a multiple of the network reduction factor. With a convolutional layer as output, the grid cells share the same weight vector. Therefore, the output layer defines a single “detector” that is applied independently to subparts of the image to construct the grided output. It ensures that the objectness score has the same meaning for a given detection unit over all the grid cells for post-prediction filtering (Sect. 2.8). It also helps the training process since all objects contribute to training the same position-agnostic set of weights regardless of their associated grid cells.
@@
An interesting side effect of having a fully convolutional structure is that we can use any input size that is a multiple of the reduction factor. Doing so will only change the size of the output grid but not the behavior of the detector at the level of a detection unit. In classical computer vision applications, the detections must be invariant to the apparent scale of an object and its resolutions. For this, the bounding boxes must be expressed in a resolution-invariant format, for example, by predicting the box sizes as fractions of the image size. The network can then be trained using various resolutions. In the classical YOLO method (starting with V2), this principle is mainly used to adjust the computational cost and detection accuracy of the model by adjusting the input resolution at prediction time.
For astronomical applications, having a resolution invariance model is not always beneficial. In our datasets, the pixel size is often valuable information regarding the instrument limitations. Due to the synthesized beam effect acting as a resolution limit, downscaling an image by a factor of two will not result in sources that look similar to sources of half the size. It goes the same way for the flux value that cannot be trivially preserved when changing the scaling of an astronomical image. For these reasons, we decided to drop the detector apparent scale invariance apparent in our method. Instead, we consider all the predicted box quantities directly related to a fixed pixel scale as constrained during training. By doing this, our detector will predict the same detection box for a specific source regardless of the input size. We could, in theory, apply our detector to the whole $32768{\times}32768$ input image of the SDC1 in a single prediction path. This approach should produce the same results as our overlapping patches split, or even better results, as it would eliminate edge effects created by the patch decomposition. However, the maximum input size remains limited by hardware constraints, and it is not always computationally efficient to use the largest possible image size for training or prediction. We use this principle in Sects 3.8 and 3.9, independently optimizing the training and prediction input sizes regarding their respective constraints.
@@ A.2 Grid resolution and detection unit density
With a fully convolutional architecture, the output grid resolution depends on the input size and the chosen network backbone reduction factor. In addition, the number of detection units per grid cell is fixed before training (Sect. 2.5). Therefore, the network structure sets the total number of detectable objects in an image in advance as a function of the input resolution, which is a common limit of most regression-based detectors. It is a substantial issue for models that are supposed to be scale and resolution-invariant, as changing the input size alters the number of output grid cells and therefore the maximum number of objects that can be detected, which contradicts their claimed resolution-invariant property. With our nonscale-invariant approach, the only quantity that matters is the density of objects at a given resolution. We consider that our pixel size represents a fixed angular scale defined by the instrument resolution limit. Increasing the input size is then equivalent to having a wider field of view, which does not change the pixel size of the objects. From there, the reduction factor and the number of detection units per grid cell can be chosen to be representative of the source density in the field.
@@
In the classical YOLO method, the multiple detection units per grid cell were not designed to handle simultaneous multiple detections. In everyday-life images, objects are considered to obstruct each other, and, in most cases, only a single object from a given scale regime can be detected by a grid cell simultaneously. In this classical formalism, the multiple detection units per grid cell are not here to allow multiple detections but to distribute the objects by apparent size or aspect ratio. Due to this construction, the default YOLO association process during the training phase was not designed to detect multiple objects of similar size centered in the same grid cell (Sect. A.3).
@@
From this, as stated in Sect. 2.5, classical detection methods answer to a high object density by lowering the reduction factor so each grid cell maps a small enough area unlikely to contain multiple objects. The issue with this approach is that it forces the network to work at a higher resolution while still reaching the necessary expressivity and having a large enough receptive field. In practice, typical 2D astronomical images with high sensitivity are so crowded that reaching the appropriate reduction factor is unrealistic, as it would result in a network that is computationally costly and very difficult to train. While using a finner output grid resolution is helpful up to some point, we also need to be capable of detecting multiple similar objects in each grid cell in order to reach the target object density. This can be achieved by switching to a prediction-aware association function as exposed in Sects. 2.6 and A.3. For our reference network backbone, we chose a reduction factor of 16 (Sect. 3.6), which is lower than the typical reduction factor of 32 of various backbones, including the YOLO-V2 architecture (darknet-19), but still high enough for our network to remain trainable and computationally efficient.
@@ A.3 Prediction-aware association
During training, the association function is responsible for finding the best two-by-two associations from two independent lists of targets and predicted boxes, all in the same grid cell (Sect. 2.6). In the classical YOLO method, the association process is based on the best theoretical match using the size priors of each detection unit. For this, it will consider that all target and predicted boxes are centered on a fictive $(0,0)$ coordinate and only compare their respective IoU, as illustrated in Fig. 23. In the classical YOLO, all detection units are implicitly supposed to represent mutually exclusive size categories, and it is considered unlikely for multiple targets of similar size to be present in the same grid cell. When it happens, the second closest prior is used, even if it is badly constrained on the target size regime.
Figure 23: Illustration of a zero-centered comparison between a target in dashed green and different size-priors in shades of red, which represents the classical YOLO association process.
To detect multiple small objects in a single grid cell, we first need multiple detection units with comparable size priors. We then use an association function based on the IoU between the targets and the current position and size of the predicted boxes. To illustrate the difference between these two approaches, we present in Fig. 24 a case study with three target boxes and four predictions from independent detection units, all in the same grid cell, and the resulting association based on both the theoretical-prior association and the best current prediction association. We observe that the first one considers that no prediction is placed correctly but that their size is relatively good. In contrast, the second one considers that the network is already good at detecting the objects and that the predicted size needs to be adjusted.
Figure 24: Comparison of the target-prediction association in a grid cell using two different association functions. The size of predicted boxes is considered close to the prior of their corresponding detection unit. The color indicates which prediction was associated with each target. The association results are illustrated for the default YOLO association function that is based on the best theoretical priors (left) and for an association function that uses the best current network prediction (right).
@@ A.4 Box matching metric
Figure 25: Comparison of three IoU-based metrics in different contexts. For all cases, the reference box is the dashed green box, and two test boxes in blue and orange are compared. The IoU-based metrics are computed between a given case and the reference.
A prediction-aware association function has to be capable of selecting which prediction is the best current representation of the target. It means we need a box-matching metric capable of discriminating between all the possible predictions in our specific context. The classical IoU matching criteria defined in Sect. 2.1 with Eq. 9 present strong limitations. To illustrate them, we compare IoU computations between a reference box and two possible test boxes in different contexts in Fig. 25. We see that the IoU is equal to zero for all couple of boxes that have no intersection regardless of their distance (Fig. 25, frame A). While it is an issue for a prediction-aware association function, the best theoretical association from the classical YOLO is not affected due to the recentering of all boxes to compare. For our specific astronomical applications where objects can be much smaller than the input size mapped by a grid cell, we need to be capable of attributing a different match score to nonintersecting boxes based on their distance to the reference. For this, we can use the Generalized IoU (GIoU, Rezatofighi et al. 2019) that makes use of the smallest box C that encloses the two boxes A and B to be compared by refining the default IoU following
@@
The GIoU can take values from -1 to 1 while still being equal to the IoU in cases where C is equal to one of the two boxes. While leveraging the first limitation, the GIoU remains insufficient to distinguish between test boxes fully contained inside the reference box. In such a context, we would like to attribute a different match score based on the distance between the box centers (Fig. 25, frame B). For this, we can use the Distance IoU (DIoU, Zheng et al. 2020) that combines the classical IoU with the ratio between the distance of the box centers $A_{c}$ and $B_{c}$ and the smallest enclosing box diagonal-length $d$ in the form
@@
where $\rho$ represents the Euclidean distance between the two center coordinates. The DIoU can take values from -1 to 1 while being equal to the IoU and GIoU in cases where the two boxes have the same center, corresponding to $\rho^{2}(A_{c},B_{c})=0$ (Fig. 25, frame C and D). This last metric is well suited for astronomical applications, but it remains limited in cases where the aspect ratio of the boxes matters. For example, all the metrics we described would give the same results for two test boxes with the same center and identical surface but different aspect ratios (Fig. 25, frame D). This could be included in a more refined matching metric like the Complete IoU (CIoU, Zheng et al. 2020).
In our YOLO-CIANNA method, all these IoU variants are only used as match criteria or as a scalar for scaling the objectness. The bounding-box-related loss remains a simple sum-of-square error on the box position and size following what was presented in Sect. 2.1 and Eqs. 6 and 7. This contrasts with what is described in the papers describing these metrics, where they are presented as alternative loss functions for object detection methods. In practice, all these matching metrics are implemented within YOLO-CIANNA and can be selected regarding the application. The predicted objectness uses the corresponding metric (rescaled between 0 and 1 if necessary) to follow Eqs. 10 and 11, so the predicted score reflects all the related subtleties.
@@ A.5 Association ordering effect
In the classical YOLO, the association is done by looping over the target list for each grid cell, which works well as they are considered unlikely to require the same detection unit. With an association process based on the current network prediction, any detection unit can be the best representation for any target regardless of its size prior. As described in Sect. 2.6, we defined our YOLO-CIANNA association function to be a search for the highest value in a matching score matrix that contains the ${\rm fIoU}$ for all the target-prediction pairs. We illustrate the effect of such an association function compared to a loop over the targets and a loop over the predictions in Fig. 26. The main advantage of our approach is that the well-detected objects will get associated first, so they will not be wrongly forced to adapt toward a different object than the one they were trying to detect. In addition, targets that are not yet well detected are more likely to be associated with a truly idling detection unit.
Figure 26: Comparison of the target-prediction association in a grid cell regarding the search order. The color indicates which prediction was associated with each target. All predictions and targets have the same size and shape, respectively. The prediction-target pairs are compared using a scaled matching metric expressed as $({\rm DIoU}+1){\times}0.5{\times}100$. The results for all pairs are displayed in a matching matrix. For each step of the association process, the rows and columns corresponding to the targets and predictions are highlighted in light red, with the intersection being the match score for this pair. For all subsequent steps, the now unavailable elements in the matrix are shown in gray. The three frames compare different association results, along with the corresponding association steps highlighted in the match score matrix, when looping over the targets (left), looping over the predictions (middle), or searching for the maximum value in the score matrix (right).
@@ A.6 Association refinements
Using a prediction-aware association implies that the network is free to distribute the diversity of objects over the detection units independently of their size prior if needed. This added degree of freedom increases training difficulty. Therefore, completely removing the theoretical box-prior association process is usually inefficient. This section describes several refinements we added to our prediction-aware association (Sect. 2.6) to constrain the accessible solutions and guide the training process. Regarding our complete loss function (Eq. 14), these refinements only affect the content of the $\mathds{1}^{\textrm{match}}$ and $\mathds{1}^{\rm void}$ masks. Most of them occur after a target-prediction couple has been selected in the association search and are verified sequentially in a mutually exclusive way, following the blue B.2 block in Fig. 5.
We emphasize that we describe all the refinements available in our method, whether or not they have been employed for the SDC1 application. Listing all the accessible refinements in the same place will allow future reference.